# Protection Degree and Migration in the Stochastic SIRS Model: A Queueing System Perspective

### v1

@@ II-A Markov Chain of Individual State
Additionally, through the transition matrix, we notice that $d_{j}(n)$ is a crucial element that affects the infected probability of a node. Intuitively, the number of infected neighbors of a node is relevant to its degree. Let the density of infected individuals at $n$ time step over the whole network be $\rho(n)$ numerically between 00 and $1$. Intuitively, we have
@@ II-B1 Markov Queueing Network of SIRS
We hereby introduce the queueing theory to an SIRS epidemic model to describe the whole process explicitly. In a propagation process, there are transitions between epidemic states and also the migration among cities. Hence, we let a place, e.g., a city or a district be an open Markov network where three epidemic states are service centers. Each epidemic state is also described as a sub-system in which, if customers are under service, they remain staying in a state. Analyzing the input and output process of each state, the transition from $S$ to $I$ can be interpreted as the removal from $S$ while the arrival at $I$. Analogously, the transition from $I$ to $R$ is the removal from $I$ and as well as the arrival at $R$. The transition from $R$ to $S$ is part of the input of $S$ and the removal from $I$, since there are mobile individuals who arrive at $S$ outside the system and a certain portion of recovered individuals will leave the system rather than return to $S$. The infected rate is denoted by $\beta$, the recovered rate is $\gamma$, the input rate is $\lambda$, the output rate is $\alpha$, and the reviving proportion is $p$. For a better understanding, the queueing system is visually illustrated in Fig. 2. We next construct the system in detail by defining its basic elements as follows.
@@
Service mechanism. All service centers $S$, $I$ and $R$ are characterized by infinite servers numerically from 00 to $\infty$. And the service time of each customer is independent identically distributed in all three systems. While the concrete time duration of service is different. The distribution service time of $S$ is taken as an exponential distribution with service rates related to the number of individuals in $I$ service center. The service time at $I$ service center is independent identically exponential distributed with the same parameter $\gamma$ of Poisson process. The service time at $R$ service center is also independent identically exponential distributed with the same parameter $\alpha$ of Poisson process.
@@
In the above mechanism, mobile people migrating to a place are all susceptible which is reasonable during the epidemic-prevent period, and the input number of people each time unit is random therefore following a Poisson distribution. While for the whole system, the output of it is all from the $R$ service center since only recovered people are allowed to leave the place. When an individual enters the system from the outside, it directly arrives at $S$ service center, then there is a probability that remains staying in $S$ or transferring to $I$ service center (be infected). An individual at $I$ has the probability to transfer to the $R$ service center. While individuals at $R$ have two choices, leaving the system or return to $S$ again. The recovery of infected nodes occurs stochastically which suits the Poisson process, therefore, the input of $R$ system is a Poisson flow. The transition from $R$ to $S$ occurs randomly and the probability of occurrence is relatively small, the Poisson process with a small rate is therefore appropriate for that flow. In terms of service mechanism, the number of individuals in an epidemic state is noticeable without limit and all coming individuals are allowed to join the service center. Hence the number of servers in each service center is endless. Susceptible individuals being infected is related to the density of infected individuals. Accordingly, the staying time in $S$ service center depends on the number of customers under $I$ service center. Individuals keep in the $I$ and $R$ service center for some time and then leave, which occurs randomly, determining the service time at both the $I$ and $R$ follows an exponential distribution.
@@ II-B2 Numerical Study on Individual Number in Stationary
Consequently, the input of $S$ can be regarded as a compound Poisson flow with the rate $\lambda+pr(t)\alpha$. Then the probability that the service is unfinished is $e^{-i_{k}(t)k{\beta}h}$. In addition, the probability that $n$ nodes come to $S$ and exactly $n-1$ nodes leave is $o(h)$. Hence, according to above calculation and Eq.11, we yield the transition probability of $S(t)$
@@
The probability that $n(n\geq 2)$ nodes leave $S$ and exactly $n-1$ nodes come is $o(h)$. Together with Eq.13, we obtain
@@
[…]
@@ III-A Transition of Epidemic State
The underlying network is a WS small-world network whose distribution of degree follows a Poisson distribution and the average path length is short, thus it fits the population network in an area. A WS small-world network is constructed by randomly reconnecting a regular network in which each node has $k$ neighbors both on the left and right sides. In our experiments, the network is initialized by $networkx$ in Python, where the scale of the network is set to be 1000, each vertex is connected to 5 neighbors on the left and right respectively, and the reconnection probability is 0.5. And we utilize $numpy.random.poisson()$ function to generate protection degrees. The experiments are carried out by setting different values 1 and 2 for the parameter of Poisson distribution and we make a comparison with the propagation process where individuals without protection degrees. The terminal time is set to be 1500 that is long enough to let the whole process be stationary. As time passes by, we record each individual’s state at each time step when the system is stationary. Eventually, we calculate the mean value of frequencies of being at I states of nodes classified by degree $k$ to evaluate the level of epidemics.
@@ III-B Queueing System of SIRS
Figure 5: The number of individuals varying with time of the three states with different parameters: The figure respectively shows the influence of each parameter on the individual number. Sub-Fig. (a) is set as $\lambda$=3, $\beta$=0.001, $\gamma$=0.7, $\alpha$=0.8 and $p$=0.995. $\alpha$, $\lambda$, $\beta$, $\gamma$ and $p$ are changed respectively from (b)-(e).
TABLE I: The results of the stationary individual number of $S$ with $\lambda$=3, $\alpha$=0.8, $p$=0.995, $<k>$=5.
@@
TABLE II: The results of the stationary individual number of $I$ with $\beta$=0.001, $\alpha=0.8$ and $<k>$=5.
@@
TABLE III: The results of the stationary individual number of $R$ with $\beta$=0.001, $\gamma=0.7$ and $<k>$=5.
@@
Figure 6: The comparison of the $S$ individual number distributions with different parameters: Five parameters influence the distribution of the individual number of $S$ differently. $\gamma$ and $\beta$ are deciding factors of $S$, where $\gamma$ is proportional to the individual number while $\beta$ works inversely. The initial setting is $\lambda$=3, $\beta$=0.001, $\gamma$=0.7, $\alpha$=0.8 and $p$=0.995. Sub-Figs. (a)-(e) display the distributions based on the original parameters changing $\alpha$, $\lambda$, $p$, $\gamma$ and $\beta$ orderly.
@@
Figure 7: The comparison of the $I$ individual number distribution with different parameters: Five parameters influence the distribution of the individual number of $S$ differently where $\lambda$ and $p$ are proportional to the individual number while $\gamma$ is inversely proportional to that. The initial setting is $\lambda$=3, $\beta$=0.001, $\gamma$=0.7, $\alpha$=0.8 and $p$=0.995. Sub-Figs.(a)-(e) display the distributions based on the original parameters changing $\alpha$, $\gamma$, $\lambda$, $\beta$ and $p$ orderly.
@@
Figure 8: The comparison of the $R$ individual number distribution with different parameters: Five parameters influence the distribution of the individual number of $S$ differently. $\lambda$ and $p$ are proportional to the individual number while $\alpha$ has an inverse effect. The initial setting is $\lambda$=3, $\beta$=0.001, $\gamma$=0.7, $\alpha$=0.8 and $p$=0.995. Sub-Figs.(a)-(e) display the distributions based on the original parameters changing $\alpha$, $\gamma$, $\lambda$, $\beta$ and $p$ orderly.
@@
The individual number varying with time is shown in Fig. 5, where plots of the number of individuals all tend to be stationary, fluctuating around a certain value after some time. In all sub-figures in Fig. 5, the blue curve denotes the number of $S$ varying with time, the red curve indicates the number of $I$, and the green one indicates the number of $R$. And we can see that in all these sub-figures, during a very short period from the beginning, the susceptible individuals grow rapidly while the infected and recovered individuals increase slowly. Then, the number of susceptible individuals decrease abruptly while at the same time, the number of individuals in infected and recovered state both increase. After $t$=1500, the number of individuals in three states becomes stationary. Fig. 5(a) displays the system with $\lambda$=3, $\beta$=0.001, $\gamma$=0.7, $\alpha$=0.8, and $p$=0.995. The individual number in three states $S$, $I$ and $R$ eventually approaches 200, 750, and 850 approximately.
@@
In addition, we illustrate the distribution of the number of individuals in $S$, $I$ and $R$ respectively given by different parameters. Different values are also set for $\alpha$, $\gamma$, $\lambda$, $\beta$ and $p$ displayed in Figs. 6, 7 and 8, where the results present an obvious normal distribution in form.
@@
Fig. 8 displays the distribution of the individual number in $R$. According to Fig. 8(a), $\alpha$=1.6 corresponds to the position of the peak on the horizontal ordinate is lower than 400, $\alpha$=0.8 corresponds to lower than 800, and the value on the horizontal ordinate of the peak with $\alpha$=0.4 get doubled approximately, matching the value between 1400 and 1600, which indicates $\alpha$ is inverse proportion to the number of $R$. In addition, the distribution with $\alpha$=0.4 has a bigger width than that with $\alpha$=0.8 and $\alpha$=1.6, whose variance is larger. On the contrary, the distribution with different $\lambda$ and $p$ is shown in Figs. 8(b) and (c) where $\lambda$=1.5, 3, 6 corresponds to the value between 250 and 500 closer to 500, the value between 750 to 1500, closer to 750 and value approaching 1750 approximately, and $p$=0.990, 0.995, 0.997 corresponds to 400, 800, 1200, both of which presenting a positive influence on the individual number of $R$. Large values of $\lambda$ and $p$ cause the distribution is on the right of the ordinate, namely, accelerating the number of individuals of $R$. Besides, the distributions in Figs. 8(d) and (e) both are overlapping, which indicates that the distribution of the individual number in $R$ is independent of $\gamma$ and $\beta$.
@@ III-C Simulation of a second wave of COVID-19 in Zhengzhou
In this subsection, we simulate the number of infected individuals in Zhengzhou, Henan province, China by utilizing the SIRS population model. According to our proposed model based on the queueing system, the infected individual number is the number of individuals in the I service center. In order to control the second wave of epidemics in Henan, the specific areas are under restriction, where people are allowed getting into the area but leaving only for essential trips. Therefore, we simulate the infected individual number in the restricted area which is regarded as an open Markov queueing network in our population model. Based on the data set of COVID-19, the second wave in Henan began on 31st July when there were domestic infected cases appearing. The restriction for the high-risk area began on 3rd August. We take the time window from 6th August to 28th August which lasts 23 days, and mainly observe the epidemics during this time window. We set the $t_{0}$=0 and the termination $T$=22. It takes time to be diagnosed, which leads to the real data delayed, while our theoretical results are timely, hence, we shift the theoretical results backward in time when comparing them.
@@
We next analyze the real situation and apply our model to real data. Most confirmed cases are in Zhengzhou, thus we take the data of Henan province as that of Zhengzhou. In the real situation of the case in Zhengzhou, most people have been vaccinated, hence, a large fraction of individuals are in the recovered state. However, the immunity got from the vaccine is not absolutely valid for the virus especially for the mutated ones. There is still the probability to get infected once contacting the infected individuals. Therefore, we regard all individuals in the area as susceptible individuals with a quite large protection degree to describe the situation that being vaccinated but not absolutely valid. Additionally, we suppose that the individuals involved are close contacts numbered 2515 which compose the underlying network of the epidemics in the high-risk area, and suppose all these individuals are in the restricted area. For the parameters in the population model, we then set appropriate values for them. In terms of the network setting, the population network is regarded as a small-world network with the total individual number $n$=2000 and the initial nearest neighbor are both one on the left and right side. We suppose that the fraction $p$ of recovered individuals being susceptible again is 0 during the second wave, then $\alpha$ is the output rate excluding the reviving flow. The input and output rate are set to be very small due to the high risk, especially for the output rate is even smaller than the input rate under restriction, and we set $\lambda$=0.01 and $\alpha$=0.005. The protection degree is set to follow a Poisson distribution with the intensity $\mu$=1 which is the expectation. The infected rate $\beta$ becomes $\beta\mu$ seven days after the first individual being infected. The recovered rate $\gamma$ is set to be 0.55.
@@ IV Conclusions and Outlook
In this paper, we study the SIRS epidemic model from two perspectives based on statistical methods. In the first model, in terms of individual, we introduce the individual protection degree considering the protective awareness in an epidemic and obtain the numerical value of stationary probability that an individual is in $S$, $I$, and $R$ state. In the second model regarding the individual number, we consider the migration of mobile people. A Markov queueing system is constructed to model the migration of people and transitions between epidemic states. Distinct from the number of individuals converging to a constant value obtained by primer studies, we achieve that the individual number in each epidemic state is convergent in probability by demonstrating the limited distributions by simulations and also obtain the expectation value. Moreover, the protection degree suppresses the epidemic level according to our model and the impact of parameters in our model, e.g., the input rate of a place and the recovered rate of infected people have a strong influence on the individual number, which inspires us to provide some suggestions for epidemic controlling theoretically.

### v2

@@ II-A Markov Chain of Individual State
Additionally, through the transition matrix, we notice that $d_{j}(n)$ is a crucial element that affects the infected probability of a node. Intuitively, the number of infected neighbors of a node is relevant to its degree. Let the density of infected individuals at $n$ time step over the whole network be $\rho(n)$ numerically between 0 and $1$. Intuitively, we have
@@ II-B1 Markov Queueing Network of SIRS
We hereby introduce the queueing theory to an SIRS epidemic model to describe the whole process explicitly. In a propagation process, there are transitions between epidemic states and also the migration among cities. Hence, we let a place, e.g., a city or a district be an open Markov network where three epidemic states are service centers. Each epidemic state is also described as a sub-system in which, if customers are under service, they remain staying in a state. Analyzing the input and output process of each state, the transition from $S$ to $I$ can be interpreted as the removal from $S$ while the arrival at $I$. Analogously, the transition from $I$ to $R$ is the removal from II and as well as the arrival at $R$. The transition from $R$ to $S$ is part of the input of $S$ and the removal from $I$, since there are mobile individuals who arrive at $S$ outside the system and a certain portion of recovered individuals will leave the system rather than return to $S$. The infected rate is denoted by $\beta$, the recovered rate is $\gamma$, the input rate is $\lambda$, the output rate is $\alpha$, and the reviving proportion is $p$. For a better understanding, the queueing system is visually illustrated in Fig. 2. We next construct the system in detail by defining its basic elements as follows.
@@
Service mechanism. All service centers $S$, II and $R$ are characterized by infinite servers numerically from 0 to $\infty$. And the service time of each customer is independent identically distributed in all three systems. While the concrete time duration of service is different. The distribution service time of $S$ is taken as an exponential distribution with service rates related to the number of individuals in $I$ service center. The service time at $I$ service center is independent identically exponential distributed with the same parameter $\gamma$ of Poisson process. The service time at $R$ service center is also independent identically exponential distributed with the same parameter $\alpha$ of Poisson process.
@@
In the above mechanism, mobile people migrating to a place are all susceptible which is reasonable during the epidemic-prevent period, and the input number of people each time unit is random therefore following a Poisson distribution. While for the whole system, the output of it is all from the $R$ service center since only recovered people are allowed to leave the place. When an individual enters the system from the outside, it directly arrives at $S$ service center, then there is a probability that remains staying in $S$ or transferring to $I$ service center (be infected). An individual at $I$ has the probability to transfer to the $R$ service center. While individuals at $R$ have two choices, leaving the system or return to $S$ again. The recovery of infected nodes occurs stochastically which suits the Poisson process, therefore, the input of $R$ system is a Poisson flow. The transition from $R$ to $S$ occurs randomly and the probability of occurrence is relatively small, the Poisson process with a small rate is therefore appropriate for that flow. In terms of service mechanism, the number of individuals in an epidemic state is noticeable without limit and all coming individuals are allowed to join the service center. Hence the number of servers in each service center is endless. Susceptible individuals being infected is related to the density of infected individuals. Accordingly, the staying time in $S$ service center depends on the number of customers under $I$ service center. Individuals keep in the II and $R$ service center for some time and then leave, which occurs randomly, determining the service time at both the II and $R$ follows an exponential distribution.
@@ II-B2 Numerical Study on Individual Number in Stationary
Consequently, the input of $S$ can be regarded as a compound Poisson flow with the rate $\lambda+pr(t)\alpha$. Then the probability that the service is unfinished is $e^{-i_{k}(t)k{\beta}h}$. In addition, the probability that $n$ nodes come to $S$ and exactly n1n-1 nodes leave is $o(h)$. Hence, according to above calculation and Eq.11, we yield the transition probability of $S(t)$
@@
The probability that $n(n\geq 2)$ nodes leave $S$ and exactly n1n-1 nodes come is $o(h)$. Together with Eq.13, we obtain
@@
$\left\{\begin{aligned} \lambda+p\alpha E[R]&=E[S]E[I]\beta k\\ E[S]E[I]\beta k&=\gamma E[I]\\ \gamma E[I]&=\alpha E[R]\end{aligned}\right.$. (18)
@@ III-A Transition of Epidemic State
The underlying network is a WS small-world network whose distribution of degree follows a Poisson distribution and the average path length is short, thus it fits the population network in an area. A WS small-world network is constructed by randomly reconnecting a regular network in which each node has $k$ neighbors both on the left and right sides. In our experiments, the network is initialized by networkxnetworkx in Python, where the scale of the network is set to be 1000, each vertex is connected to 5 neighbors on the left and right respectively, and the reconnection probability is 0.5. And we utilize $numpy.random.poisson()$ function to generate protection degrees. The experiments are carried out by setting different values 1 and 2 for the parameter of Poisson distribution and we make a comparison with the propagation process where individuals without protection degrees. The terminal time is set to be 1500 that is long enough to let the whole process be stationary. As time passes by, we record each individual’s state at each time step when the system is stationary. Eventually, we calculate the mean value of frequencies of being at I states of nodes classified by degree $k$ to evaluate the level of epidemics.
@@ III-B Queueing System of SIRS
Figure 5: The number of individuals varying with time of the three states with different parameters: The figure respectively shows the influence of each parameter on the individual number. Sub-Fig. (a) is set as $\lambda$=3, $\beta$=0.001, $\gamma$=0.7, $\alpha$=0.8 and pp=0.995. $\alpha$, $\lambda$, $\beta$, $\gamma$ and $p$ are changed respectively from (b)-(e).
TABLE I: The results of the stationary individual number of $S$ with $\lambda$=3, $\alpha$=0.8, pp=0.995, <k><k>=5.
@@
TABLE II: The results of the stationary individual number of II with $\beta$=0.001, $\alpha=0.8$ and <k><k>=5.
@@
TABLE III: The results of the stationary individual number of $R$ with $\beta$=0.001, $\gamma=0.7$ and <k><k>=5.
@@
Figure 6: The comparison of the $S$ individual number distributions with different parameters: Five parameters influence the distribution of the individual number of $S$ differently. $\gamma$ and $\beta$ are deciding factors of $S$, where $\gamma$ is proportional to the individual number while $\beta$ works inversely. The initial setting is $\lambda$=3, $\beta$=0.001, $\gamma$=0.7, $\alpha$=0.8 and pp=0.995. Sub-Figs. (a)-(e) display the distributions based on the original parameters changing $\alpha$, $\lambda$, $p$, $\gamma$ and $\beta$ orderly.
@@
Figure 7: The comparison of the $I$ individual number distribution with different parameters: Five parameters influence the distribution of the individual number of $S$ differently where $\lambda$ and $p$ are proportional to the individual number while $\gamma$ is inversely proportional to that. The initial setting is $\lambda$=3, $\beta$=0.001, $\gamma$=0.7, $\alpha$=0.8 and pp=0.995. Sub-Figs.(a)-(e) display the distributions based on the original parameters changing $\alpha$, $\gamma$, $\lambda$, $\beta$ and $p$ orderly.
@@
Figure 8: The comparison of the $R$ individual number distribution with different parameters: Five parameters influence the distribution of the individual number of $S$ differently. $\lambda$ and $p$ are proportional to the individual number while $\alpha$ has an inverse effect. The initial setting is $\lambda$=3, $\beta$=0.001, $\gamma$=0.7, $\alpha$=0.8 and pp=0.995. Sub-Figs.(a)-(e) display the distributions based on the original parameters changing $\alpha$, $\gamma$, $\lambda$, $\beta$ and $p$ orderly.
@@
The individual number varying with time is shown in Fig. 5, where plots of the number of individuals all tend to be stationary, fluctuating around a certain value after some time. In all sub-figures in Fig. 5, the blue curve denotes the number of $S$ varying with time, the red curve indicates the number of $I$, and the green one indicates the number of $R$. And we can see that in all these sub-figures, during a very short period from the beginning, the susceptible individuals grow rapidly while the infected and recovered individuals increase slowly. Then, the number of susceptible individuals decrease abruptly while at the same time, the number of individuals in infected and recovered state both increase. After tt=1500, the number of individuals in three states becomes stationary. Fig. 5(a) displays the system with $\lambda$=3, $\beta$=0.001, $\gamma$=0.7, $\alpha$=0.8, and pp=0.995. The individual number in three states $S$, II and $R$ eventually approaches 200, 750, and 850 approximately.
@@
In addition, we illustrate the distribution of the number of individuals in $S$, II and $R$ respectively given by different parameters. Different values are also set for $\alpha$, $\gamma$, $\lambda$, $\beta$ and $p$ displayed in Figs. 6, 7 and 8, where the results present an obvious normal distribution in form.
@@
Fig. 8 displays the distribution of the individual number in $R$. According to Fig. 8(a), $\alpha$=1.6 corresponds to the position of the peak on the horizontal ordinate is lower than 400, $\alpha$=0.8 corresponds to lower than 800, and the value on the horizontal ordinate of the peak with $\alpha$=0.4 get doubled approximately, matching the value between 1400 and 1600, which indicates $\alpha$ is inverse proportion to the number of $R$. In addition, the distribution with $\alpha$=0.4 has a bigger width than that with $\alpha$=0.8 and $\alpha$=1.6, whose variance is larger. On the contrary, the distribution with different $\lambda$ and $p$ is shown in Figs. 8(b) and (c) where $\lambda$=1.5, 3, 6 corresponds to the value between 250 and 500 closer to 500, the value between 750 to 1500, closer to 750 and value approaching 1750 approximately, and pp=0.990, 0.995, 0.997 corresponds to 400, 800, 1200, both of which presenting a positive influence on the individual number of $R$. Large values of $\lambda$ and $p$ cause the distribution is on the right of the ordinate, namely, accelerating the number of individuals of $R$. Besides, the distributions in Figs. 8(d) and (e) both are overlapping, which indicates that the distribution of the individual number in $R$ is independent of $\gamma$ and $\beta$.
@@ III-C Simulation of a second wave of COVID-19 in Zhengzhou
In this subsection, we simulate the number of infected individuals in Zhengzhou, Henan province, China by utilizing the SIRS population model. According to our proposed model based on the queueing system, the infected individual number is the number of individuals in the I service center. In order to control the second wave of epidemics in Henan, the specific areas are under restriction, where people are allowed getting into the area but leaving only for essential trips. Therefore, we simulate the infected individual number in the restricted area which is regarded as an open Markov queueing network in our population model. Based on the data set of COVID-19, the second wave in Henan began on 31st July when there were domestic infected cases appearing. The restriction for the high-risk area began on 3rd August. We take the time window from 6th August to 28th August which lasts 23 days, and mainly observe the epidemics during this time window. We set the $t_{0}$=0 and the termination TT=22. It takes time to be diagnosed, which leads to the real data delayed, while our theoretical results are timely, hence, we shift the theoretical results backward in time when comparing them.
@@
We next analyze the real situation and apply our model to real data. Most confirmed cases are in Zhengzhou, thus we take the data of Henan province as that of Zhengzhou. In the real situation of the case in Zhengzhou, most people have been vaccinated, hence, a large fraction of individuals are in the recovered state. However, the immunity got from the vaccine is not absolutely valid for the virus especially for the mutated ones. There is still the probability to get infected once contacting the infected individuals. Therefore, we regard all individuals in the area as susceptible individuals with a quite large protection degree to describe the situation that being vaccinated but not absolutely valid. Additionally, we suppose that the individuals involved are close contacts numbered 2515 which compose the underlying network of the epidemics in the high-risk area, and suppose all these individuals are in the restricted area. For the parameters in the population model, we then set appropriate values for them. In terms of the network setting, the population network is regarded as a small-world network with the total individual number nn=2000 and the initial nearest neighbor are both one on the left and right side. We suppose that the fraction $p$ of recovered individuals being susceptible again is 0 during the second wave, then $\alpha$ is the output rate excluding the reviving flow. The input and output rate are set to be very small due to the high risk, especially for the output rate is even smaller than the input rate under restriction, and we set $\lambda$=0.01 and $\alpha$=0.005. The protection degree is set to follow a Poisson distribution with the intensity $\mu$=1 which is the expectation. The infected rate $\beta$ becomes $\beta\mu$ seven days after the first individual being infected. The recovered rate $\gamma$ is set to be 0.55.
@@ IV Conclusions and Outlook
In this paper, we study the SIRS epidemic model from two perspectives based on statistical methods. In the first model, in terms of individual, we introduce the individual protection degree considering the protective awareness in an epidemic and obtain the numerical value of stationary probability that an individual is in $S$, II, and $R$ state. In the second model regarding the individual number, we consider the migration of mobile people. A Markov queueing system is constructed to model the migration of people and transitions between epidemic states. Distinct from the number of individuals converging to a constant value obtained by primer studies, we achieve that the individual number in each epidemic state is convergent in probability by demonstrating the limited distributions by simulations and also obtain the expectation value. Moreover, the protection degree suppresses the epidemic level according to our model and the impact of parameters in our model, e.g., the input rate of a place and the recovered rate of infected people have a strong influence on the individual number, which inspires us to provide some suggestions for epidemic controlling theoretically.